# Koopmon trajectories in nonadiabatic quantum-classical dynamics

### 21 Dec 2023 v1

@@ Acknowledgments.
We are all greatly indebted to Darryl Holm, who is part of the collaboration working on the bohmion scheme and provided extremely valuable advice during our discussion meetings with him at Imperial College London. CT also wishes to thank Robert MacKay for his encouragement to work in this direction and for his insistence on applying the new model to problems in physics and chemistry. It turned out very rewarding advice. We also thank all the other people who supported our efforts and provided further motivation during this work, especially Victor S. Batista, Michael Berry, Denys Bondar, Irene Burghardt, Basile Curchod, Eberhard Gross, Caroline Lasser, Robert Littlejohn, Giovanni Manfredi, Phil Morrison, Micheline B. Soley, and John Tully. This work was made possible through the support of Grant 62210 from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. FGB was partially supported by a start-up grant from the Nanyang Technological University.

### 17 Jun 2024 v2

@@ Acknowledgments.
We are all greatly indebted to Darryl Holm, who is part of the collaboration working on the bohmion scheme and provided extremely valuable advice during our discussion meetings with him at Imperial College London. CT also wishes to thank Robert MacKay for his encouragement to work in this direction and for his insistence on the need to apply the new model to problems in physics and chemistry. It turned out very rewarding advice. We also thank all the other people who supported our efforts and provided further motivation during this work, especially Victor S. Batista, Michael Berry, Denys Bondar, Irene Burghardt, Basile Curchod, Eberhard Gross, Caroline Lasser, Robert Littlejohn, Giovanni Manfredi, Phil Morrison, Micheline B. Soley, John Tully and Tomasz Tyranowski. We would also like to thank Delyan Zhelyazov for his recent work on the adaptive-timestep implementation. Finally, we thank the reviewers for pointing out the importance of nonadiabatic couplings and for drawing our attention to the literature on the Koopman operator. This work was made possible through the support of Grant 62210 from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. Financial support by the Leverhulme Research Project Grant RPG-2023-078 is also greatly acknowledged. FGB was partially supported by a start-up grant from the Nanyang Technological University.
Appendix A Role of the nonadiabatic coupling
This Appendix expands on the phenomenology associated to the nonadiabatic processes observed in the Tully models from Section 4.
In the context of the Tully models, the koopmon trajectories model the interaction between nuclear degrees of freedom (classical subsystem), described by the classical trajectories $(q_{a},p_{a})$, and electronic degrees of freedom (quantum subsystem), described by the density matrices $\hat{\varrho}_{a}$. In the adiabatic regime, the nuclear dynamics take place on the two-level potential energy surfaces (PESs), which are shown in the top row of Figure 15. These surfaces are sometimes referred to as adiabatic PESs, reflecting the fact that they are given by the eigenvalues of the electronic Hamiltonian ($\widehat{H}_{\operatorname{el}}(q)={H_{0}(q){\boldsymbol{1}}+H_{1}(q)\hat{\sigma}_{x}+H_{3}(q)\hat{\sigma}_{z}}$), with the eigenvectors forming the so-called adiabatic basis.
Figure 15: PESs (top) and nonadiabatic couplings (bottom) for the Tully models.
In nonadiabatic dynamics, however, one is usually not only interested in the nuclear dynamics on adiabatic PESs, but also in the exchange of quantum amplitudes between the surfaces. The typical picture involves the Born–Huang expansion
$\Psi(r,t)=\psi_{1}(r,t)v_{1}(r)+\psi_{2}(r,t)v_{2}(r)$
of the molecular wave function $\Psi(\cdot,t)\in L^{2}(\mathbb{R},\mathbb{C}^{2})$ in the adiabatic basis $\{v_{1}(r),v_{2}(r)\}\subset\mathbb{C}^{2}$ made available by the spectral problem for the electronic Hamiltonian. In this context, the so-called nonadiabatic couplings (NACs) are often used to identify regions in the PES landscape where nonadiabatic processes are likely to be observed. In regions where the PESs come close to each other, so-called nonadiabatic transitions occur, and the different Tully models were designed to create a platform for testing models to capture these effects.
NACs are expressed in terms of NAC vectors, which are particularly important for computing the standard jump probability for the fewest switches surface hopping algorithm introduced by Tully [74]. For two-level systems defined on a one-dimensional coordinate space, the NAC vector (in this case a real number) is given by
$d^{1,2}(r):=\langle v_{1}(r)\mid\nabla v_{2}(r)\rangle=\frac{\langle v_{1}(r)\mid\nabla H_{\operatorname{el}}(r)v_{2}(r)\rangle}{\lambda_{1}(r)-\lambda_{2}(r)},$ (23)
where $\lambda_{j}(r),\,j\in\{1,2\}$, denotes the corresponding eigenvalues (and hence PES) of the adiabatic basis vectors $v_{j}(r)$ at a given position $r\in\mathbb{R}$. The last term in (23) provides an alternative way to compute the NAC, and the proof of the second equality goes back to [12].
Figure 15 shows a plot of the NACs (blue) for the three Tully models. As we can see, the values of $d^{1,2}(r)$ are indeed large in regions where the PESs (top) are close to each other. Furthermore, formula (23) reveals that the NAC vectors have a singularity for positions with $\lambda_{1}(r)=\lambda_{2}(r)$, and it is known that the adiabatic approximation breaks down near these points.
As we can see in Figures 2, 4 and 6, regions of strong nonadiabatic coupling cause significant deformations of the initial Gaussian wavepacket entering from the left. Moreover, the presence of negative values in these regions indicates purely quantum effects. Once these negativities disappear, the population levels reach a constant value, indicating that there is no further exchange of quantum amplitudes. For the Tully models I and II, we find that in regions with strong nonadiabatic coupling there is a certain loss of accuracy in reproducing the correct population levels as predicted by the fully quantum code. However, the koopmons reproduce the correct population levels before entering and after leaving the regions of strong nonadiabatic coupling. For Tully III, the plot of the NAC shows that the nonadiabatic coupling reaches its maximum around $x\approx-6$. The koopmons are unique in qualitatively capturing the correct population dynamics for all times (see Figure 7). Both the Ehrenfest and bohmion methods fail to capture the reflection (no negative momentum for the trajectories), while the koopmons reproduce negative momenta as predicted by the fully quantum simulation. Since only the reflected trajectories revisit the region of strong nonadiabatic coupling, more transitions between the PESs take place between times $t=2000$ and $t=3000$ only for the koopmons, thereby increasing the population on the ground state.